OpenStudy (anonymous):

~~Fanning ALL who helps plus Medal~~ Which polynomial is a perfect square trinomial? 49x^2 − 28x + 16 4a^2 − 20a + 25 25b^2 − 20b − 16 16x^2 − 24x − 9

OpenStudy (anonymous):

second equation

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

Okay Thank you and the reason why I questioned you is because I looked up the answer and one person said it was A the other said it was D Loll thanks

OpenStudy (anonymous):

Hey want to know how to find out?

OpenStudy (anonymous):

Sure why not I used to know this so well about 5 months ago

OpenStudy (anonymous):

\[ 49x^2 − 28x + 16 \] Take the middle term \(-28\) and then divide by two times the first term \(2\times 49\), then square it:\[ \left(\frac{-28}{2\times 49}\right)^2=\left(-\frac{2}{7}\right)^2=\frac {4}{14} \]This must be equatl to the last term. Since \(16\neq 4/14\), this one isn't a perfect square.

OpenStudy (anonymous):

ahuh

OpenStudy (anonymous):

Wait that formula is a bit wrong

OpenStudy (anonymous):

Okay go ahead

OpenStudy (anonymous):

You have to multiply the first term back i

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

so which one is the right answer?

OpenStudy (anonymous):

A much more intuitive approach is to try to complete the square.

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

How'd they get d? in the link you showed me just now?

OpenStudy (anonymous):

They said \[ d=\frac{b}{2a} \]

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

You want \(e=0\) to have a perfect square which means :\[ c = \frac{b^2}{4a} \]By the way \[ \frac{b^2}{4a}=a\left(\frac{b}{2a}\right)^2 \]

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Okay let's try it again ok? \[ 49x^2 − 28x + 16 \]So \[ \frac{b^2}{4a}=\frac{(-28)^2}{4(49)} = 4\neq 16 \]

OpenStudy (anonymous):

Now with \[ 4a^2 − 20a + 25 \]We use \[ \frac{b^2}{4a}=\frac{(-20)^2}{4(4)}=\frac{400}{16}=25=c \]So it is a complete square.

OpenStudy (anonymous):

I'm listening

OpenStudy (anonymous):

Do you understand how I put b and a into the equation?

OpenStudy (anonymous):

hm... what i did was take a times sqrt of c and see if that's the same number as b

OpenStudy (anonymous):

I understand and I see penguin

OpenStudy (anonymous):

yes yes yes...

OpenStudy (anonymous):

Loll

OpenStudy (anonymous):

You did \[ a\sqrt c = b \]? penguin?

OpenStudy (anonymous):

yeah, but negative

OpenStudy (anonymous):

Well, \[ b^2=4ac \]Would work, or even \[ b=\sqrt{4ac} \]But not \[ a\sqrt c = b \]...

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

@wio so the aswer is a?

OpenStudy (anonymous):

The second one happens to be correct.

OpenStudy (anonymous):

Ok thank you both